You charge an initially uncharged 89.9-mF capacitor through a 35.7-? resistor by means of a 9.00-V battery having negligible internal resistance. Find the time constant of the circuit. What is the charge of the capacitor 1.09 time constants after the circuit is closed?
Answer #1

t = RC

= 3.21 *10^-3 Seconds


Q= CV (1-e(-t/RC)) = 0.809*0.6637 = 0.537 mC
Answer #2

Time constanrt = R*C = 35.7*89.9*10^-3 = 3.2 s


Q = Qo*(1 - e^(-t/T))


= C*Vo*(1 - e^(-1.09))


= 89.9*10^-3*9*(1 - e^(-1.09))


= 5.37*10^-4 C